OpenStudy (anonymous):

Please help me and show me how you got the answer! A cone has a radius of 8 inches and a height of 12 inches. What is the exact volume of the cone? A. 32π in3 B. 96π in3 C. 256π in3 D. 768π in3

jimthompson5910 (jim_thompson5910):

Use the formula V = (1/3)*pi*r^2*h V = volume r = radius h = height

OpenStudy (anonymous):

I am having trouble doing that

jimthompson5910 (jim_thompson5910):

In this case, r = 8 h = 12

OpenStudy (anonymous):

8*12 ?

jimthompson5910 (jim_thompson5910):

V = (1/3)*pi*r^2*h V = (1/3)*pi*8^2*12 V = ???

OpenStudy (anonymous):

I don't understand

jimthompson5910 (jim_thompson5910):

what is 8^2

OpenStudy (anonymous):

64

jimthompson5910 (jim_thompson5910):

then multiply that with 1/3 and 12

OpenStudy (anonymous):

768

jimthompson5910 (jim_thompson5910):

that's too large

OpenStudy (anonymous):

oh.. I did the wrong math... 96??

jimthompson5910 (jim_thompson5910):

64*12 = ???

OpenStudy (anonymous):

= 768

jimthompson5910 (jim_thompson5910):

now you multiply that by 1/3 ie divide by 3

OpenStudy (anonymous):

256

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

good, so the final answer is \[\large 256\pi\] you're welcome